I am very confused can someone help me with ths quadratic problem:
x^2+x-6
Thanks alot

Answer by stanbon(75887)

x^2+x-6
(x+3)(x-2) is the answer if you are only factoring.
--------------
If you have an equation you proceed as follows:
x^2+x-6=0
(x+3)(x-2)=0
x = -3 or x=2
===========================
Cheers,
Stan H.